A vessel of volume 0.2m3 contained N2 at 1.013 bar. 15oC. If 0.2kg of N2 is now pumped into the vessel, calculate the new pressure when the vessel has returned to its initial temperature. The molecular weight of N2 is 28kg/kmol.


1
Expert's answer
2021-02-05T04:54:54-0500

"V = 0.2 m^3 = 200 L"

P = 1.013 bar = 0.99975 atm

T = 15 C = 288 K

Moles of Nitrogen, n = "\frac{PV}{RT}"

= 0.99975 atm "\times" 200 L / (0.0821 L-atm/mol.K "\times" 288 K)

                                = 8.46 moles

Moles of nitrogen added, n = mass/ molar mass

                                            ="\frac{ 0.2 *\times10^3 g} { 28 g/mol}"

                                          = 7.14 moles

Total moles = 8.46 + 7.14 = 15.6 moles

New pressure P = n R T / V

                         = "\frac{15.6 moles \times 0.0821 L-atm/ mol.K \times 288 K } {200 L}"

                         = 1.84 atm = 1.864 bar